OpenStudy (anonymous):

Bridgette brings d dollars to an amusement park. The park charges an admission fee of $18. Each game at the park costs $2.50. Which inequality shows the number of games, g, which Bridgette can play?

OpenStudy (anonymous):

how many dollars does she bring?

OpenStudy (anonymous):

If Bridgette plays g games, 1. pay one time admission 18, 2 g*2.5 for the games she played 18+g*2.5<=d

OpenStudy (anonymous):

g*2.5<=d-18 g<=(d-18)/2.5

OpenStudy (anonymous):

Why <=, because your expenses should be less than or equal to the amount you have in pocket. As you cannot play on credit :)

OpenStudy (anonymous):

I put 18 + 2.5g <=d and got it wrong. That is what brought me here. Think I should question this answer???

OpenStudy (anonymous):

try g<=(d-18)/2.5. It is same as the earlier, but in different format. If it also got wrong, definitely, some thing messy with question . :)